AGW Commissioning and LUCIFER Commissioning on 13-14 September 2008

Observers: J. Hill, N. Ageorges, W. Seifert, M. Pedani and numerous other LUCIFERians on mtn
More Observers: D. Miller and A. Rakich in Tucson
Software Support: M. Delapena in Tucson
Telescope Operators: D. Gonzalez Huerta and W. Wack

Summary

Comment heard in German at 6AM in the control room: This was a real night of instrument commissioning! We lost the first 1.5 hours after sunset to changing the DSP board in Jet. We had no Jet-MCSPU problems for the rest of the night. We also had no problems with the chillers and the control room was a comfortable temperature. Things went quite smoothly with only the occasional hang and restarting of software. The telescope was in better temperature equilibrium than last night, but the seeing was worse. LUCIFER images were in the 0.7 to 1.2 arcsec FWHM range - decent images but really ruining the averages started last week with 0.3 arcsec.

We were able to reliably track, guide and do active optics through LUCIFER presets. But a variety of little problems make it the case that it typically takes 2 or 3 tries before a preset succeeds. These problems are gradually being understood and corrected. Doug and Andrew were able to take off-Y-axis WFS data to confirm the sign of the correction. The correct sign should be included in tomorrow's patch of GCS.

The biggest telescope problem that we don't understand are the 0.3-0.5 arcsec jumps in the images that we first noticed in the IRTC data on Jupiter.

The only thing that ruined the night was that Nancy had a bad headache. Kai and Peter took her down to the emergency room in Safford at 4AM in case it was altitude sickness. As of 6:20AM, we are still waiting a report from the doctor.

Nighttime Activities

01:20 Open Shutter Doors with anticipation of sky flats, but problems with MCSPU again even though it worked fine all afternoon

01:30 John and David replacing the DSP card in MCSPU with phone assistance by Dan C.. We have not yet replaced the system (CPU) card, although we have the spare in hand.

02:55 Close shutter doors.

03:00 To zenith to work on LUCIFER electronics

03:22 Re-open Shutter Doors - T=8degC D=1 degC. Mostly clear.

03:25 Problems with AGw; oacserverlog hung, won't start correctly. agw-control is slow to give a login prompt, but rebooting agw-control didn't help this problem. We deduce that PMAC may be hung since oacserver stops after printing some motor-control type commands. Finally power cycling AGw solves problem.

03:30 Resetting telescope main axis encoders.

04:03 Pointto target below horizon (John needs more sleep!) and preset fails correctly.

04:05 Pointto M5_1539 TRACK - MCSPU made a slewing error.

04:08 readGuideCam has long delay before giving me an image. Had to netconfig kill GCSL, and even that wouldn't stop it.

04:15 After restarting GCSL, found star at offset -30 30. Center star with IE=-115 CA=-35 (was this the wrong star?, yes)

04:25 pointto M5_1539 ACTIVE (Why did the focus jump when we did this?)

04:29 pointto M5_1539 ACTIVE grabs the star. Now we are asking WFS to fix fairly strong coma.

04:35 pointto M5_1539 ACTIVE gs=5 PA=INDEF > 213.59

# Target coordinates for source M5_1539
  1. :38:31.376 21:12:04.40 RADEC_SKY 2000.0 J2000 0.0 0.0 5.0 R 0.0 H_K 1.4
    # Guide Star coordinates for GS05
  2. :38:22.372 21:08:54.79 RADEC_SKY 2000.0 J2000 0.0 0.0 11.84 R 0.33 B_V 0.63

04:45 Preset with LUCI, didn't find star was 1.5 arcmin North. Now IE=-20 CA=-53

05:07 Preset with LUCI gs=offset=60 arcsec, star appears dead center in acquisition image

05:11 Preset with LUCI offset=120 arcsec, star at 230, 264 image 269

05:15 Preset with LUCI offset=240 arcsec star at 190,247 image 270

Why did this experiment succeed tonight when it failed last night?

05:24 Close vent doors - Wind is ~8 m/sec on roof.

05:25 LUCIFER preset, finds star well centered (278,263) at x=-1.5 Y=540. Doug reports 100 nm rms after two corrections.

05:31 Blocking guide corrections with PCS, and blocking WFS with Doug.

Images are jumping around in guide images and in the LUCIFER image with 0.3-0.5 arcsec amplitude. It is not caused by guide corrections. Vibration? Atmospheric turbulence? Wind shake? This is likely the same problem as in Issue 1482 with IRTC images.

05:38 Manual collimation, recenter with guiding

05:50 LUCIFER preset with guide star 3 arcmin off-axis x=-2.5 y=504.2 (star at 168, 239)

05:54 Sent 4 manual triggerWFS (last sent image is 667). Starting with 668, we are not sending the corrections. End at 675.

06:10 Resend preset followed by 6 triggerWFS corrections; first correction had 300 nm astigmatism, images 676-681

06:17 Start WFS with no send to primary 682-685

06:25 Preset found star at (162,244) with probe x=-2.4 y=504.2 = Send 3 triggerWFS starting at 686, first correction of 280 nm rms to <200 nm

06:30 Guide and ActOpt loop on-axis starting at #690-698, star found (221,250)

06:42 Guide mode off-axis for focus sequence

Net Z4 corrections (relative to nominal +29 mm AGw focus): 0 nm, +130 nm, +260 nm, -260 nm. -130 nm, 0 nm, +260 nm, and then I lost track when he repeated. Scale is 165 nm of Z4 per mm of Gregorian focus motion.

06:53 Active mode on-axis. 20-sec exposures, 120 nm rms.

29 mm, 33 mm, 25 mm, 33 mm, 29 mm, (but it really went to 0 mm, because we did "setfocus 29" rather than "setfocus -p 29".

07:10 Stopped SX Mirror Ventilation for ~1 minute to verify that it was the source of the image jumping. It wasn't.

General conclusion is that the AGw focus value of 29 mm is correct for LUCIFER, and we should repeat the Z4 experiment in better seeing for filling in the first decimal place.

07:18 Preset on-axis ACTIVE to a high elevation field, star went through guide box, but then away

07:22 Resend preset, star oscillated but locked at second try. Az velocity is -170 arcsec/sec.

Guiding is having some oscillation problems here at high elevation - not enough time for the azimuth axis to respond.

GCS needs to allow slightly more time for the offset of the star onto the hotspot whenever the star is near the edge of the acquistion box, or the target is at high elevation.

07:34 Block guiding at PCS

Funny problems with GCSGUI and info from 107. Maybe this was because Michele was making event database queries on lbtmu107.

07:42 Preset in TRACK mode - Sep 14 00:42:02 lbtmu105 LBT_IIF: [I] guide loop didn't stop within timeout.

07:51 Preset to a different field, but still near EL=87. (96,384) This one also oscillates at 0.4 Hz. Guiding looks stable at 0.1 Hz.

07:57 Preset to another field at lower elevation. Walter is checking image quality in all the LUCIFER filters.

08:07 Preset in TRACK mode:
Sep 14 01:07:21 lbtmu105 LBT_IIF: [I] left PMC Mirror GoToPositionAbs Command Failed because Chase state not IDLE

Sep 14 01:07:21 lbtmu107 LBT_left_GCS: PROFILING: WFS: Zernikes sent...

01:09:40 irs hangs in between two LUCIFER presets. LUCIFER saw the success at 01:08:10, but not the success at 01:09:32 MST. We send Michele to restart irs. The process was there, but hung.

08:24 Another LUCIFER preset.

08:32 LUCIFER preset ACTIVE, star at 61,142 but guider didn't pick it up.

08:39 GCSL seems hung (along with GCSGUI). Wouldn't respond to netconfig stop. Restarted GCSL and GCSGUI. Sending the same 01:32 preset again

Walter says narrowband K images are 0.7 arcsec. Doug was estimating 1.5 arcsec images from the guider. Telescope elevation is ~40 degrees. Seeing monitor is reporting 2-2.4 arcsec.

08:47-09:14 /tmp/wfscimage000760 - 794 Off-axis guiding and ActOpt. It ran happliy for half an hour.

09:24 onaxis presetTelescope to star +40 in elevation (~36 to 73). Huge amount of focus. Should be able to do this correction by reading ActOpt PSF buffers. Change focus by hand and then let ActOpt loop do its magic.

02:36 Walter sends off-axis preset to same star. Both GCSL and irs hang and had to be restarted. Unclear if the two hangs are related. After restarting the off-axis guidestar appears at 395,33.

10:24 Start guiding and ActOpt off-axis, and off the Y-axis

Doug and Andrew Measurements of Off-Y-axis wavefront correction

10:45 Changed signs on delta and phi, may have some cross talk or may be field corrections

wfsc 5 no extra aberration

wfsc 6 z5 = 1500 with -field_phi, +rotation_delta, both z5 and z6 change

wfsc 7 with +field_phi, -rotation_delta, almost all z5

wfsc 8 remove z5 (-1500) loose 1000 z5

10:50 BS9188, GUIDE, gstar=16: no star found, or very out of focus

10:55 BS9188, TRACK, PA=322: out of focus on-axis star

Recentered the pointing zeropoints. IE=-30 CA=-60

And we remembered that changing authorization rereads the pointing model and thus cancels your zeropoint corrections. Apparently this is also the reason for the big change in focus when the collimation model global offsets are read.

11:00 Nancy has a severe headache. Kai and Peter take her down the mountain to Safford hospital in case the headache is related to altitude sickness.

11:03 BS9188, GUIDE, gstar=16: acquire guide star

wfsc 9 94 -57 -37 66

11:06 BS9188, GUIDE, gstar=16 PA=270

wfsc 10 107 -25 -47 60

wfsc 11 107 -27 -47 60

11:03 BS9188, GUIDE, gstar=16: acquire guide star

wfsc 12 90 -63 -35 68

LUCIFER running dither script test

11:34 Kill GCS, although it may not have been guilty in this instance

11:36 Preset LUCIFER Again with closer off-axis star

11:42 irs hung mid-preset

12:02 - 12:04 IIF has level 5 event that says preset completed at 05:0245 MST. But two minutes later, irs has not returned anything to LUCIFER. Michele restarts irs and we preset again.

12:08 GCS dies

12:20 This is the best example time for the LUCIFER offsets and Issue 1633.

12:24 Telescope stopped tracking because HBS shutdown. Apparently this is a new HBS error and not one of the routine old ones. Lost 10 minutes.

12:35 At zenith for sky flats

13:05 Close dome. T=8degC D=-9degC! Wind 4m/sec Sky Photometric

-- JohnHill - 14 Sep 2008
Topic revision: r10 - 14 Sep 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback